OpenStudy (anonymous):

hi can anyone tell me how to find the matrix whose nullspace is((2 1 0),(3 0 1))

OpenStudy (fwizbang):

Since (210) and (301) are in the null space, then A(210)^T = A(301)^T =0 which means the rows of A are orthogonal to both of those vectors. So A= (1 -2 -3 2 - 4 -6 3 -6 -9) works.(There are an infinite set of possibilities, A needn't even be 3 by 3.)